# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

### v1

@@ Appendix A Proof for Theorem 2
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=$ $\int_{\{F_{\mu}(x)\leq 1/2\}}\left(x-\mathbb{E}[X]+\frac{\theta^{*}_{2}}{\sqrt{2\pi}}\right)^{2}d\mu(x)$
$+$ $\int_{\{F_{\mu}(x)>1/2\}}\left(x-\mathbb{E}[X]-\theta^{*}_{2}(\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}})\right)^{2}d\mu(x)$
$=$ $\int_{\{F_{\mu}(x)\leq 1/2\}}(x-\mathbb{E}[X])^{2}d\mu(x)$
$+$ $\sqrt{\frac{2}{\pi}}\theta^{*}_{2}\int_{\{F_{\mu}(x)\leq 1/2\}}(x-\mathbb{E}[X])d\mu(x)$
$+$ $\frac{(\theta^{*}_{2})^{2}}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$+$ $\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])^{2}d\mu(x)$
- $2\theta^{*}_{2}\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}}\Big{)}d\mu(x)$
Observe that
$=$ $\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}$
- $\frac{1}{\sqrt{2\pi}}1_{\{F_{\mu}(x)>1/2\}}\Big{)}d\mu(x)$
$=$ $\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}\Big{)}d\mu(x)$
$+$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}(x-\mathbb{E}[X])1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$=$ $\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}\right)$
$+$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}(x-\mathbb{E}[X])1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x),$ (86)
where the last equality follows from (80). Similarly,
$=$ $\int_{\mathbb{R}}\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}\Big{)}^{2}d\mu(x)$
- $\frac{1}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=$ $\operatorname{Var}(X)-2\theta^{*}_{2}\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(x)>1/2\}}\right)$
$+(\theta^{*}_{2})^{2}\frac{\pi-1}{2\pi}$
where the last equality is due to the definition of $\theta^{*}_{2}$ in (83). ∎
Next, we goes to $\theta_{2}\leq 0$. Under the condition $\theta_{2}\leq 0$, we deduce from (29) and (12) that
@@
Indeed, for the case $F_{\mu}(x)\leq 1/2$, by a calculation similar to (32), but with $\theta_{2}\leq 0$, we get $\Psi\big{(}\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}\big{)}=1-F_{\mu}(x)$, which implies
$\begin{split}&t^{\Theta}(x)=\theta_{1}+\theta_{2}\Psi^{-1}(1-F_{\mu}(x))\\ =&\theta_{1}+\theta_{2}\Phi^{-1}(1-F_{\mu}(x))=\theta_{1}-\theta_{2}\Phi^{-1}(F_{\mu}(x)),\end{split}$
@@
$\begin{split}&\min_{\theta_{1}\in\mathbb{R},\eta\geq 0}J(\theta_{1},\eta):=\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &+\int_{\{F_{\mu}(x)\leq 1/2\}}\Big{(}\theta_{1}+\eta\Phi^{-1}(F_{\mu}(x))-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &\hbox{subject to}\quad g(\theta_{1},\eta):=-\eta\leq 0.\end{split}$ (89)
@@
$\mathbb{E}[X]\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]$
$+$ $\eta\text{Var}\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)$
- $\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]-\lambda/2=0.$ (90)
@@
$\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]$
$=$ $\mathbb{E}[Z1_{\{\Phi(Z)\leq 1/2\}}]=\mathbb{E}[Z1_{\{Z\leq 0\}}]=-\frac{1}{\sqrt{2\pi}},$ (91)
$\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)\leq 1/2\}}\right]$
$=$ $\mathbb{E}[Z^{2}1_{\{\Phi(Z)\leq 1/2\}}]=\mathbb{E}[Z^{2}1_{\{Z\leq 0\}}]=\frac{1}{2}.$ (92)
@@
Consider $f(x):=\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)\leq 1/2\}}$ and set $x^{*}:=\sup\{x\in\mathbb{R}:F_{\mu}(x)\leq 1/2\}$. As $F_{\mu}$ is nondecreasing and $\Phi^{-1}$ is strictly increasing, for $x<x^{*}$, $f(x)=\Phi^{-1}(F_{\mu}(x))$ is nondecreasing with $f(x)\leq\Phi^{-1}(1/2)=0$; for $x\geq x^{*}$, we simply have $f\equiv 0$. Since $f$ is nondecreasing on $\mathbb{R}$, an argument similar to (40) implies $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)\geq 0$. The result then follows by repeating the same arguments in the proof of Le,,a 1. Specifically, we deal with the two cases where $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)$ is strictly positive and zero separately. In each case, we solve (93) (in place of (38) in Lemma 1) and find that $H_{J}+\lambda H_{g}$ now takes the form
@@
Lemma 12.
Moreover, the minimum value is
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$ (96)
$=$ $\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)^{2}.$
Proof.
Since (44) is equivalent to (89) by taking $\theta_{2}=-\eta$, Lemma 11 directly provides the unique minimizer $\theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})$, given as in (95), for (44). By repeating all the calculations in Lemma 10 (with the two sets $\{F_{\mu}(x)\leq 1/2\}$ and $\{F_{\mu}(x)>1/2\}$ switched and $\theta_{2}$ replaced by $\eta^{*}$ in (94)), we obtain the minimum value formula. ∎
Finally we can get our theorem. If (13) holds, the value in (84) is smaller or equal to that in (96). Hence, the minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should be the one associated with the value (84), which is given by (83). On the other hand, if (13) fails, the value in (96) is smaller than that in (84). The minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should then be the one associated with the value (96), which is given by (95).
Appendix B Proof of Corollary 1
@@
where $\nu$ is the distribution of generator input, a Gaussian vector. The reason is by transform of random variables, the optimal solution the inner discriminator problem of (1) can be transformed to be the optimal solution of inner discriminator problem in (99). Let $\pi^{*}(x,y,\Theta)$ be the solution of the inner discriminator problem, then from Danskin’s theorem [32, Proposition B.25], its gradient over $\Theta$ is
@@
Under W2 with $d=1$, the gradient over vector $\theta=(\theta_{1},\theta_{2})$
@@
$=$ $-2\;[\mathbb{E}_{\mu}[X]-(\theta_{1}+\theta_{2}\mathbb{E}[Z])\;,$ (101)
$\mathbb{E}_{\pi*}[XZ]-(\theta_{1}\mathbb{E}[Z]+\theta_{2}\mathbb{E}[Z^{2}])]^{T},$
since from matrix derivative
@@
which results (16) by letting (B) equals to $0,0]^{T}$.
For (17) we swapping the role of $X$ and $G_{\Theta}(Z)$ in Lemma 3, and from (B)
@@
Note that the optimal joint distribution $\pi*$ will equal to that of $(t^{\Theta}(Y),Y)$. From (104) and $t^{\Theta}(Y)=X$ we reach (103).
Appendix C Detailed Proof for Lemma 1
@@
$\Bigg{|}$ $\Big{(}\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\Big{)}^{\frac{1}{q}}$-
$\Big{(}\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\Big{)}^{\frac{1}{q}}\Bigg{|}\leq f^{\mu}(d)$ (105)
@@
where the marginals with respect to the first and second variables are given by $\mu$ and $\nu$ respectively. The case for $q=2$ is proved by [17, Corollary 3]. Now for $q=1$,
$\int_{\omega\in\Omega_{G}}W_{1}(\omega^{T}X,\hat{X}_{G})d\omega\leq\int_{\omega\in\Omega_{G}}W_{2}(\omega^{T}X,\hat{X}_{G})d\omega$
$\leq$ $\bigg{(}\int_{\omega\in\Omega_{G}}W^{2}_{2}(\omega^{T}X,\hat{X}_{G})d\omega\bigg{)}^{1/2}\leq f^{\mu}(d)$
where the first inequality is from $W_{1}\leq W_{2}$ by the definition of Wasserstein distances (see [33, Remark 6.6]) and the second one is from Jensen’s inequality. Then from triangle inequality, given $\omega$
$|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|$
$\leq$ $W_{q}(\omega^{T}X,\hat{X}_{G}),q=1,2,$
@@
$\int_{\omega\in\Omega_{G}}|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|^{q}d\omega$
$\leq(f^{\mu}(d))^{q}.$ (106)
Finally, by triangle inequality
$\Bigg{|}$ $\left(\int_{\omega\in\Omega_{G}}W^{q}_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))\right)^{1/q}$
$-\left(\int_{\omega\in\Omega_{G}}W^{q}_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))\right)^{1/q}\Bigg{|}$
@@
Note that when $q=1$, one do not need triangle inequality to make the upper-bound valid. Then with (106), we have (105).
@@
which equals to solving
@@
Finally, it is easy to see that (23) with $q=2$ equals to (49). Here we focus on the term
@@
$=\frac{1}{\Gamma(1/2)U_{w}}\int_{0}^{\infty}\bigg{(}\frac{y}{U_{w}}\bigg{)}^{1/2-1}e^{-y}dy$
$=\frac{1}{\Gamma(1/2)}\int_{0}^{\infty}z^{-1/2}e^{-zU_{w}}dz,$
where the second line follows from the change of variable $z=y/U_{w}$. Note that the above calculation is motivated by [34, Example 3.2b.2]. It follows that (111) equals to
@@
which equals to (50) from $\mathbb{E}_{U_{w}}[U_{w}\exp(-zU_{w})]=-\frac{\partial}{\partial z}M_{U_{w}}(-z)$ where we recall $M_{U_{w}}(z):=\mathbb{E}[e^{zU_{w}}]$ is the moment generating function of $U_{w}$. Moreover, from [34, Theorem 3.2a.1], $M_{U_{w}}(z)$ has a closed form as
@@
Appendix D Extension of Lemma 1 for $q=1$
To get a solution as (24) for $\Theta$ under $q=1$, we sacrifice the approximation error in Lemma 1 by modifying (105) (106) to get upcoming (116). Details are as follows. Given $\omega$ and $\Theta$, from triangle inequality
$|W_{1}(\omega^{T}X,\omega^{T}\Theta Z)-W_{1}(\hat{X}_{G},\hat{Z}_{G})|$
$\leq$ $W_{1}(\omega^{T}X,\hat{X}_{G})+W_{1}(\omega^{T}\Theta Z,\hat{Z}_{G}),$
where $\hat{Z}_{G}\sim\mathcal{N}(0,\tilde{\sigma}^{2}_{\Theta})$ is independent of $\omega$ with
$\tilde{\sigma}^{2}_{\Theta}=\frac{\mathbb{E}[\|\Theta Z\|^{2}]}{d}=\frac{\mathrm{Tr}(\Theta\Theta^{T})}{d}.$
Then (106) is modified as
$\int_{\omega\in\Omega_{G}}|W_{1}(\omega^{T}X,\omega^{T}\Theta Z)-W_{1}(\hat{X}_{G},\hat{Z}_{G})|d\omega$
$\leq$ $f^{\mu}(d)+f^{\Theta Z}(d),$ (113)
where $f^{\Theta Z}(d)$ is similarly defined as (48) by replacing $\mu$ with $\Theta Z$ and still $O(d^{-1/8})$. Then for every $\Theta$, following previous steps to prove (105) from (106) we have
$\Bigg{|}$ $\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|]d\omega$ (114)
$-\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\hat{Z}_{G}|]d\omega\Bigg{|}$
is bounded by $O(d^{-1/8})$ for large $d$. Following the proof of reaching (108), now we aim to solve
$\min_{\Theta}|\tilde{\sigma}_{x}-\sqrt{\mathrm{Tr}(\Theta\Theta^{T})/d}|].$ (115)
and $\Theta$ meeting
$\tilde{\sigma}^{2}_{x}\mathbf{I}=\Theta\Theta^{T}$ (116)
will be the solution.

### v2

@@ Appendix A Proof for Proposition 1
$\begin{split}\eta^{*}&=\frac{\operatorname{Cov}\left(X,\Psi^{-1}_{0}(F_{\mu}(X))\right)}{\operatorname{Var}(-h(Z))}=-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\\ \theta^{*}_{1}&=\mathbb{E}[X]-\eta^{*}\mathbb{E}\left[-h(Z)\right]=\mathbb{E}[X]-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\mathbb{E}[h(Z)],\end{split}$
Now we can prove Proposition 1. By the linearity of covariances,
$\begin{split}&\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)+\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)=\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))+\Psi^{-1}(1-F_{\mu}(X))\right).\end{split}$
If (9) holds, the above implies that the value in (92) is smaller than or equal to that in (94). Hence, the minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should be the one associated with the value (92), which is given by (10). On the other hand, if (9) fails, the same argument shows that the value in (94) is smaller than that in (92). The minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should then be the one associated with the value (94), which is given by (93).
Appendix B Proof of Lemma 6
Under the condition $\theta_{2}\leq 0$, we deduce from (30) and (12) that
@@ Appendix B Proof of Lemma 6
Indeed, for the case $F_{\mu}(x)\leq 1/2$, with $\theta_{2}\leq 0$, we get
$\begin{split}\theta_{1}+\theta_{2}\Phi^{-1}(1-F_{\mu}(x))=\theta_{1}-\theta_{2}\Phi^{-1}(F_{\mu}(x)),\end{split}$
@@
$\begin{split}&\min_{\theta_{1}\in\mathbb{R},\eta\geq 0}J(\theta_{1},\eta):=\int_{\{F_{\mu}(x)>1/2\}}\Big(\theta_{1}-x\Big)^{2}F_{\mu}^{\prime}(x)dx+\int_{\{F_{\mu}(x)\leq 1/2\}}\Big(\theta_{1}+\eta\Phi^{-1}(F_{\mu}(x))-x\Big)^{2}F_{\mu}^{\prime}(x)dx\\ &\hbox{subject to}\quad g(\theta_{1},\eta):=-\eta\leq 0.\end{split}$ (97)
@@
$\mathbb{E}[X]\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]+\eta\text{Var}\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)$
$-\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]-\lambda/2=0.$ (98)
@@
$\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]=\mathbb{E}[Z1_{\{\Phi(Z)\leq 1/2\}}]=\mathbb{E}[Z1_{\{Z\leq 0\}}]=-\frac{1}{\sqrt{2\pi}},$ (99)
$\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)\leq 1/2\}}\right]=\mathbb{E}[Z^{2}1_{\{\Phi(Z)\leq 1/2\}}]=\mathbb{E}[Z^{2}1_{\{Z\leq 0\}}]=\frac{1}{2}.$ (100)
@@
Consider $f(x):=\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)\leq 1/2\}}$ and set $x^{*}:=\sup\{x\in\mathbb{R}:F_{\mu}(x)\leq 1/2\}$. As $F_{\mu}$ is nondecreasing and $\Phi^{-1}$ is strictly increasing, for $x<x^{*}$, $f(x)=\Phi^{-1}(F_{\mu}(x))$ is nondecreasing with $f(x)\leq\Phi^{-1}(1/2)=0$; for $x\geq x^{*}$, we simply have $f\equiv 0$. Since $f$ is nondecreasing on $\mathbb{R}$, an argument similar to (50) implies $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)\geq 0$. Specifically, we deal with the two cases where $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)$ is strictly positive and zero separately. In each case, we solve (101) (in place of (37)) and find that $H_{J}+\lambda H_{g}$ now takes the form
@@
By repeating all the calculations in Lemma 5 (with the two sets $\{F_{\mu}(x)\leq 1/2\}$ and $\{F_{\mu}(x)>1/2\}$ switched and $\theta_{2}$ replaced by $\eta^{*}$ in (102)), we obtain the minimum value formula.
Appendix C Proof of Corollary 1
@@
where $\nu$ is the distribution of the generator input, a Gaussian vector. The reason is that, by transforming random variables, the optimal solution of the inner discriminator problem of (1) can be transformed to be the optimal solution of the inner discriminator problem in (105). Let $\pi^{*}(x,y,\Theta)$ be the solution of the inner discriminator problem, then from Danskin’s theorem [5, Proposition B.25], its gradient over $\Theta$ is
@@
Under $q=2$ with $d=1$, the gradient over vector $\theta=(\theta_{1},\theta_{2})$
@@
$=$ $-2\;[\mathbb{E}_{\mu}[X]-(\theta_{1}+\theta_{2}\mathbb{E}[Z])\;,\mathbb{E}_{\pi*}[XZ]-(\theta_{1}\mathbb{E}[Z]+\theta_{2}\mathbb{E}[Z^{2}])]^{T},$ (107)
since the matrix derivative
@@
which results (17) by letting (107) equals to $0,0]^{T}$.
For (18) we swapping the role of $X$ and $G_{\Theta}(Z)$ in Lemma 3, and from (107)
@@
Note that the optimal joint distribution $\pi*$ will equal that of $(t^{\Theta}(Y),Y)$. From (110) and $t^{\Theta}(Y)=X$ we reach (109).
Appendix D Detailed Proof for Lemma 1
@@
$\Bigg|$ $\Big(\int_{\omega\in\Omega_{G}}\inf_{\pi_{\omega}}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\Big)^{\frac{1}{q}}-\Big(\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\Big)^{\frac{1}{q}}\Bigg|\leq f^{\mu}(d)$ (111)
@@
where the marginals with respect to the first and second variables are given by $\mu$ and $\nu$ respectively. The case for $q=2$ is proved by [28, Corollary 3]. Now for $q=1$,
$\int_{\omega\in\Omega_{G}}W_{1}(\omega^{T}X,\hat{X}_{G})d\omega\leq\int_{\omega\in\Omega_{G}}W_{2}(\omega^{T}X,\hat{X}_{G})d\omega\leq\bigg(\int_{\omega\in\Omega_{G}}W^{2}_{2}(\omega^{T}X,\hat{X}_{G})d\omega\bigg)^{1/2}\leq f^{\mu}(d)$
where the first inequality is from $W_{1}\leq W_{2}$ by the definition of Wasserstein distances (see [34, Remark 6.6]) and the second one is from Jensen’s inequality. Then from triangle inequality, given $\omega$
$|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|\leq W_{q}(\omega^{T}X,\hat{X}_{G}),q=1,2,$
@@
$\int_{\omega\in\Omega_{G}}|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|^{q}d\omega\leq(f^{\mu}(d))^{q}.$ (112)
Finally, by the triangle inequality
$\Bigg|$ $\left(\int_{\omega\in\Omega_{G}}W^{q}_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))\right)^{1/q}-\left(\int_{\omega\in\Omega_{G}}W^{q}_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))\right)^{1/q}\Bigg|$
@@
Note that when $q=1$, one does not need the triangle inequality to make the upper-bound valid. Then with (112), we have (111).
@@
which equals solving
@@
Finally, it is easy to see that (24) with $q=2$ equals (60). Here, we focus on the term
@@
$=\frac{1}{\Gamma(1/2)U_{w}}\int_{0}^{\infty}\bigg(\frac{y}{U_{w}}\bigg)^{1/2-1}e^{-y}dy=\frac{1}{\Gamma(1/2)}\int_{0}^{\infty}z^{-1/2}e^{-zU_{w}}dz,$
where the second line follows from the change of variable $z=y/U_{w}$. Note that the above calculation is motivated by [23, Example 3.2b.2]. It follows that (117) equals to
@@
which equals to (61) from $\mathbb{E}_{U_{w}}[U_{w}\exp(-zU_{w})]=-\frac{\partial}{\partial z}M_{U_{w}}(-z)$ where we recall $M_{U_{w}}(z):=\mathbb{E}[e^{zU_{w}}]$ is the moment generating function of $U_{w}$. Moreover, from [23, Theorem 3.2a.1], $M_{U_{w}}(z)$ has a closed form as